Test the hypothesis using the P-value approach. Ho: p 0.55 versus H1 p0.55 n-150, X-78, α-0.10 Perform the test using the P-value approach. P-value(Round to four decimal places as needed.)Test the hypothesis using the P-value approach. Be sure to verify the requirements of the test. Ho: p=0.72 versus H1 : p#0.72 n 500, x-350, o 0.05 Is npo (1-Po) 210? Select the correct choice below and fill in the answer box to complete your choice. (Type an integer or a decimal. Do not round.) OA. No, because npo (1-Po) - B. Yes, because nPo(1-po,-.Some have argued that throwing darts at the stock pages to decide which companies to invest in could be a successful stock-picking strategy. Suppose a researcher decides to test this theory and randomly chooses 200 companies to invest in. After 1 year, 106 of the companies were considered winners; that is, they outperformed other companies in the same investment class. To assess whether the dart-picking strategy resulted in a majority of winners, the researcher tested Hop 0.5 versus Hp>0.5 and obtained a P-value of 0.1981. Explain what this P-value means and write a conclusion for the researcher. (Assume α is 0.1 or less.) Choose the correct explanation below. O A. About 106 in 200 samples will give a sample proportion as high or higher than the one obtained if the population proportion really is greater than 0.5 O B. About 106 in 200 samples will give a sample proportion as high or higher than the one obtained if the population proportion really is 0.5. O C. About 20 in 100 samples will give a sample proportion as high or higher than the one obtained if the population proportion really is greater than 0.5 0 D. About 20 in 100 samples will give a sample proportion as high or higher than the one obtained if the population proportion really is 0.5In a clinical trial 24 out of 868 patients taking a prescription drug daily complained of flulike symptoms. Suppose that it is known that 2.2% of patients taking competing drugs complain offl like symptoms. Is there sufficient evidence to conclude that more than 2.2% of this drugs users experience flulike symptoms as a side effect at the α = 0.01 level of significance? Because neo (1-Po)-D,10, the sample size is[ hypothesissatisfied. Round to one decimal place as needed.) I 5% of the population size, and the sample V the requirements for testing the is given to be random, cannot be reasonably assumed to be random, is given to not be random, can be reasonably assumed to be randomA certain drug can be used to reduce the acid produced by the body and heal damage to the esophagus due to acid reflux The manufacturer of the drug claims that more than 89% o patients taking the drug are healed within 8 weeks in clinical trials, 206 of 227 patients suffering from acid reflux disease were healed after 8 weeks. Test the manufacturers claim at the α= 0.01 level of significance. Because npo (1-Po,-D ,10, the sample size isL , 5% of the population size, and the sample the requirements for testing the V satisfied. hypothesis (Round to one decimal place as needed.) can be reasonably assumed to be random is given to be random, cannot be reasonably assumed to be random is given to not be random,
Answer #1

Ans:

1)

sample proportion=78/150=0.52

Test statistic:

z=(0.52-0.55)/sqrt(0.55*(1-0.55)/150)

z=-0.739

P(z<-0.739)=0.2301*

*(if z is rounded off to -0.74,p-value will be=0.2296)

2)Option B is correct.

3)np0(1-p0)=868*0.022*(1-0.022)=18.7>10

population size not given,sample is not given random

4)np0(1-p0)=500*0.72*(1-0.72)=100.8

Yes,because 500*0.72*(1-0.72)=100.8